OpenStudy (anonymous):

I am to use the Gauss Jordan Method to solve for the system of equations. If the system has infinitely many solutions I am to give the z arbitrary. I have the following equation I am to solve for which is going to be x=5y+4z=1 and 3x-2y+3z=-2 for these I have arrived at the answers of -7z-1/13, 9z-5/13, z Have I worked these correctly?

hero (hero):

x + 5y + 4z = 1 3x + 2y + 3z = -2 Guess a value of z such as z = 1 Then x + 5y + 4(1) = 1 3x + 2y + 3(1) = -2 x + 5y + 4 = 1 3x + 2y + 3 = -2 x + 5y = -3 3x + 2y = -5 Now you have a system in two equations x and y which you should be able to solve.

OpenStudy (anonymous):

Okay when I work it this time I have \[\frac{-7z-12 }{ 13 }, \frac{ 9z-5 }{ 13}, z \]

OpenStudy (anonymous):

Still not getting the right answer here. Any help?

OpenStudy (fibonaccichick666):

If I recall, gauss jordan involves reducing the coefficient matrix right?

OpenStudy (anonymous):

So z is the free variable?

OpenStudy (fibonaccichick666):

Can you write the coefficient matrix? Then can you write it in reduced row echelon?

OpenStudy (fibonaccichick666):

oh and wio, yea z is free (arbitrary)

OpenStudy (anonymous):

\[ \begin{bmatrix} 1&5&4\\ 3&-2&3 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ -2 \end{bmatrix} \]

OpenStudy (anonymous):

Hmm, @ShortStaticBurst I'll show how my work and we'll see if your answer matches up.

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

I'm doing \(R_2 = R_2-3R_1\) \[ \begin{bmatrix} 1&5&4\\ 0&-17&-9 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ -5 \end{bmatrix} \]

OpenStudy (anonymous):

I'm doing \(R_2 =5 R_2+17R_1\) \[ \begin{bmatrix} 1&5&4\\ 0&0&23 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ -8 \end{bmatrix} \]

OpenStudy (anonymous):

actually, doing it this way makes \(y\) the free variable.

OpenStudy (anonymous):

But it doesn't matter.

OpenStudy (anonymous):

Now I am subtracting \(5y\) in the top equation \[ \begin{bmatrix} 1&4\\ 0&23 \end{bmatrix} \begin{bmatrix} x\\ z \end{bmatrix} = \begin{bmatrix} 1-5y\\ -8 \end{bmatrix} \]

OpenStudy (anonymous):

Actually, let me double check my work for a moment...

OpenStudy (anonymous):

@ShortStaticBurst Did you use a different method than what I am doing?

OpenStudy (anonymous):

I think my method is a bit... confusing. Let me consider an easier route.

OpenStudy (anonymous):

Yes I did. But I don't trust the method the book suggests.

OpenStudy (anonymous):

We add a new equation: \(z=z\), and it gives us:\[ \begin{bmatrix} 1&5&4\\ 3&-2&3\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ -2\\ z \end{bmatrix} \]Does this make sense to you?

OpenStudy (anonymous):

I'm doing \(R_2 = R_2-3R_1\) \[ \begin{bmatrix} 1&5&4\\ 0&-17&-9\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ -5\\ z \end{bmatrix} \]

OpenStudy (anonymous):

I used the augmented matrix from the book. two equations expressing x and y in terms of z

OpenStudy (anonymous):

There are many ways to approach this problem, so I don't know which method your book did in particular.

OpenStudy (anonymous):

I'm doing \(R_2 = -R_2/17\) \[ \begin{bmatrix} 1&5&4\\ 0&1&9/17\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\ y\\ z \end{bmatrix} = \begin{bmatrix} 1\\ 5/17\\ z \end{bmatrix} \]

OpenStudy (anonymous):

At this point, I can back substitute.

OpenStudy (anonymous):

I have: \(y + 9z/17 = 5/17\) So this means \[y = \frac{5-9z}{17}\]

OpenStudy (anonymous):

Then we have in the top row: \[ x+\frac{5-9z}{17}+4z=1 \]

OpenStudy (anonymous):

Whoops, that should be: \[ x+5\left(\frac{5-9z}{17}\right)+4z=1 \]

OpenStudy (anonymous):

The question is confusing because there are infinite solutions as far as I can tell. In fact one of the three choices to answer with has that it could be a solution set where z is any real number

OpenStudy (anonymous):

Yeah, the solution will be a line.

OpenStudy (anonymous):

There is 1 free variable.

OpenStudy (anonymous):

I really find the whole thing horribly confusing. It seems to be you should solve it with three spaces for the solution set or two spaces and a z for the solution set. This question is not one I have found to be easy to come by.

OpenStudy (anonymous):

All you are doing is adding a third equation.. it could be \(z=t\):\[ x=5y+4z=1 \\ 3x-2y+3z=-2 \\ z=t\]or if could be \(y=t\), or even \(x=t\).

OpenStudy (anonymous):

In this case \(t\) can be any number. It could be 2, 0, or -122. It's just representing a value that \(z\), \(x\), or \(y\) might be set to...

OpenStudy (anonymous):

Consider if you have just the equation \(x+y=2\). It's obvious this is a line.\[ y = -x+2 \]

OpenStudy (anonymous):

Since there is one equation and two variables, there is one free variable.

OpenStudy (anonymous):

Okay.